OpenStudy (anonymous):

Numbers in standard form (a+bi) (1-3i)/(2+i), the answer I got was (-2-7i), taking it to SF is -2/5-7i/5 is that the standard form or did I screw up?

OpenStudy (anonymous):

@Jack1 if you could take a quick look ~(;W;)~

OpenStudy (jack1):

hey minato ;D divided by complex numbers (ur *i* here is sqrt -1, yeah? )

OpenStudy (jack1):

remember that 1/i = - i ...yeah?

OpenStudy (anonymous):

Yush i=sqrt(-1)

OpenStudy (anonymous):

Did not remember that ono

OpenStudy (jack1):

so to work it out, i find it's easier to times top and bottom to remove the i so\[\frac {(1-3i)}{(2+i)} \rightarrow \frac {(1-3i) \times (2-i)}{(2+i) \times (2-i)}\] so focus on the denominators here, pretty little trick: FOIL it out like normal first: (2+i) x (2-i) = 2x2 + 2 (-i) + 2i×2 +i(-i) = 4 - (i^2) and i^2 = -1 so = 4 - (-1) = 5 so denominator turns to 5 so equation is now: \[\frac {(1-3i) \times (2-i)}{5}\] can u foil out the top line and see how u go from here dude?

OpenStudy (jack1):

sorry, doubled up there, should be *(2+i) x (2-i) = 2x2 + 2 x (-i) + 2 x i +i(-i) [not 2i x 2)

OpenStudy (anonymous):

(1-3i)(2-i) = 2-i-6i+4i^2 = 2-7i+4i^2= 2-4-7i=-2-7i i^=-1

OpenStudy (anonymous):

-2-7i over 5? Does that mean I was right?

OpenStudy (jack1):

do over: (1- 3i ) (2 - i) = 1x2 + -i + -6i + 3i^2 = 2 -i -6i + 3(-1) = 2 - 3 -7i => so = -1 - 7i

OpenStudy (jack1):

so that's numerator and 5 is denominator u can either fraction it up or decimal? up to u

OpenStudy (anonymous):

God I am such a fish stick, I kept writing -3i (-i)=4i^2

OpenStudy (jack1):

haha, all good it too me 2 attempts 2, i forgot the 3 in the first go ;D

OpenStudy (anonymous):

Sanka you Jack1-san, again. *0*

OpenStudy (jack1):

all good dude, anytime hey